A Pesticide applicator needs to have 100 gallons of a pesticide solution with 8% copper. He has several barrels in his shop containing copper solutions; one is a 2% solution and the other is a 9% solution. How many gallons of each will be required to obtain 100 gallons of an 8% copper solution?
Answer by jorel555(1290)

.09n+.02(100-n)=.08(100)
.09n+2-.02n=8
.07n=6
n=6/.07=85.71 gallons of 9% copper;100-85.71=14.29 gallons of 2% solution!!!!!!!!!!!